# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ Abstract
I analyzed the risk-coping strategies among factory worker households in early 20th-century Tokyo. I digitized and analyzed a unique daily longitudinal household budget survey to determine how consumption was impacted by idiosyncratic shocks. I found that while the households were so vulnerable that the shocks impacted their consumption levels, the estimated income elasticity for food consumption was relatively low in the short run. The analysis using adverse health shock confirms the robustness of the results. The result of mechanism analysis suggests that credit purchases with local retailers smoothed short-run food consumption. Despite the potential loss of profit, retailers in a competitive situation allowed consumers to trade on credit. This shows the roles of informal credit institutions in mitigating vulnerability among urban worker households.
@@ 1 Introduction
Economic studies that analyze the relationship between historical economic development and financial institutions have focused on the alleviation of constraints on corporate financing through official financial systems (Acemoglu et al. 2005; Levine 2005). However, systematic research on the association between household financial institutions and consumption smoothing during industrialization remains limited. How did working-class households respond to idiosyncratic shocks and smooth their consumption before public insurance was established? Development economics research has confirmed that consumption smoothing is an essential mechanism for mitigating poverty and explaining human capital formation. [^1] Evidence shows that markets in contemporary developing countries are often imperfect. Thus, the households are vulnerable to risk, and idiosyncratic shocks impact the levels of their consumption. The early representative literature includes, for example, Townsend (1994), Townsend (1995), and Ravallion and Chaudhuri (1997). Morduch (1999), Townsend (1995), and Dercon (2004) provide reviews of the earlier studies. Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions. Since shocks are not fully insured in the markets, smoothness of consumption is also an important factor in human capital formation. Foster (1995) investigates the shock of the 1988 floods in Bangladesh and shows that children’s growth patterns are influenced by the inefficiency of the credit markets. Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall events in early life increased the survival probabilities of girls in rural India. Gertler and Gruber (2002) also shows that consumption smoothing against health shocks is imperfect in Indonesia. Surprisingly, however, economic history studies have not yet provided an answer to this important question. We still know little about how historical households maintained their livelihoods during industrialization.
This study assesses risk-coping behavior among the working-class households in early twentieth-century Tokyo. I digitized and analyzed a unique daily longitudinal budget survey on factory worker households conducted in a representative manufacturing area. The full insurance hypothesis is strongly rejected for overall consumption expenditure, indicating that factory worker households in the historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks. However, expenses on indispensable consumption categories were partially smoothed in the short run. Credit purchases in local retailers played an important role in mitigating adverse idiosyncratic shocks, whereas the other risk-mitigation strategies were impractical. Credit purchases did not require an explicit contract and thus often led to a loss in sales for the retailers. However, to remain competitive, retailers allowed credit transactions so that their customers remained regular. This shows that informal credit institutions in urban local communities functioned as informal insurance networks in rural economies.
This study’s primary contribution is to add analytical perspectives on consumption smoothing behavior to economic history studies. Studies of how working-class households compensated for income loss during the late 19th to the early 20th century have focused on Britain. These studies used cross-sectional variation in income to estimate the elasticities of the temporary income sources. Their detailed analyses have revealed the roles of financial institutions and labor supply adjustments in compensating for income loss in working-class households (e.g., Scott and Walker 2012). [^2] They found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractual savings, and insurance. The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use savings institutions. Horrell and Oxley (2000) highlight the important role of labor supply adjustments among family members to manage the income losses of the head in working-class households in late 19th-century Britain. Kiesling (1996) and Boyer (1997) focused on the risk-mitigation strategies among the cotton workers households in Lancashire during the recession caused by the Civil War. However, there is little evidence of how those risk-coping strategies smoothed their consumption. [^3] Other strands of literature provide supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005), building image on the historical consumers in the long-run (Francks 2009), and describe the relationship between consumption and production (Gordon 2013). Hence, this study provides a rare window into urban household financial behavior during industrialization by utilizing unique household longitudinal budget data from prewar Japan. It also implements a systematic panel data analysis to offer comparable estimates of the income elasticity for future studies.
The second contribution is revealing the roles of informal credit institutions in historical metropolis. Evidence from development economics studies indicates that the availability of informal credit institutions relates to the stability of consumption in rural economies, especially when the availability of formal financial institutions is limited. Udry (1994) showed that state-contingent loans among relatives and neighbors could partially insure against idiosyncratic shocks on rural households in northern Nigeria. A detailed field survey revealed that asymmetric information was no longer important in those private credit contracts (Udry 1990). [^4] Fafchamps and Lund (2003) also presented evidence that the households in the rural Philippines used informal networks with relatives and friends as a risk-mitigation strategy for idiosyncratic shocks. Related studies consider risk-sharing among households with similar cultural and institutional backgrounds. Grimard (1997) showed that insurance among households in the same ethnic group could partially mitigate idiosyncratic shocks in the multiethnic economy of Côte d’Ivoire. Munshi and Rosenzweig (2016) presented the importance of rural insurance networks among workers in the same caste in the context of spatial wage inequality in India. Nonetheless, there is little evidence on the function of informal credit institutions in the historical context. Ogasawara (2024) observed the role of official financial institutions, such as postal savings and pawnshops, in compensating temporary income losses in the early 20th-century Osaka. [^5] Note that Osaka was middle scale city of approximately half the size of Tokyo City circa 1920 (Online Appendix Table A.1). However, the function of informal credit institutions in consumption smoothing has not been considered due to limited data availability. By considering the roles of both formal and informal financial institutions simultaneously, this current study documents the contribution of informal credit transactions by local retailers.
@@
Relatedly, an examination of how, and the extent to which, the local society could ensure household consumption provides valuable insights into existing urban history literature. After World War I, working-class households in Tokyo could establish and sustain their families (Nakagawa 1985). The results of this study suggest that the expansion of urbanization and credit transactions through the war underpins this traditional historical view. Another suggestion relates to the findings by Munshi and Rosenzweig (2016), who linked spatial wage inequality and men’s mobility in India with the availability of risk-mitigation networks among workers in the same caste. From this perspective, local credit institutions with lower barriers may motivate rural workers to urbanize, which may partially explain why Tokyo grew so fast circa WWI. This hypothesis may be worthy of testing in future research.
The third contribution is to offer a case study of a metropolis in a developing economy. Extant development economics studies on consumption smoothing behavior have focused particularly on the rural economies (e.g., Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002). [^6] Recent field experimental studies also focus on village economies (Comola and Prina 2023; Macours et al. 2022). While Townsend (1995), Skoufias and Quisumbing (2005), and Aguila et al. (2017) provided case studies on middle-sized contemporary economies, there is little evidence on agglomeration economies like prewar Tokyo. Urbanization is, however, now a worldwide phenomenon. The proportions of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020). This requires a deeper understanding of consumption behavior among urban working households in the economic development process. Moreover, there is little research on the role of financial institutions among the household sector in agglomeration economies. This study considers a set of homogeneous households in a representative manufacturing area in prewar Tokyo, which provides a valuable empirical setting for testing risk-sharing strategies in the historical metropolis. [^7] Focusing on a sample that shares similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). Caution is needed when comparing results, as the modern financial system differs in many ways from the past. However, my findings show that informal credit institutions in the local urban community could mitigate vulnerability among households similar to rural developing economies (Udry 1994; Fafchamps and Lund 2003).
@@ 2.1 Japanese Economy circa the First World War
Between 1885 and 1915, the Japanese real GNP grew annually by 2.6%, whereas the population increased by 1.1%. [^8] Faster economic than population growth implies that modern economic growth had started in this period as defined by Simon Kuznets (Miyamoto 2008, pp. 60; 64). In 1915, the percentage of the employed population in the primary, secondary, and tertiary industries were 63%, 20%, and 18%, respectively. Based on domestic genuine production, these industries accounted for just under 30%, over 30%, and approximately 40%, respectively. Soon after the outbreak of WWI, exports from European countries declined. Japan increased exports to the Asian market on behalf of those European countries and exported munitions and foodstuffs to the European Allies, leading to a current account surplus. The import substitution expanded the domestic market and led to a rise in the machinery, metal, and chemical sectors. A worldwide shortage of ships also stimulated the development of shipbuilding, which proliferated. The development of the shipbuilding industry further stimulated demand for the machinery and steel industries. Through this process of heavy and chemical industrialization, the total number of enterprises increased. [^9] Descriptions about the process of heavy chemical industrialization are based on Nakamura (1971, p. 128), Nakamura and Odaka (1989, p. 10), and Sawai and Tanimoto (2016, p. 250). Year-on-year growth in the number of incorporated firms was 17.8%, 22.3%, and 13.5% in 1918, 1919, and 1920, respectively. These are the top three of the 34 years from after the Russo-Japanese War to 1940 (Nakamura and Odaka 1989, p. 22). Figure 1 shows that the GNP growth from 1915–19 was 11.5% in nominal terms (4.4% in real terms), the highest growth throughout the 1910s and 1920s.
This rapid industrialization impacted the Japanese economy in many ways. Due to export-led economic growth, large inflows of gold increased domestic currency balances, causing prices to soar. Although wages began to rise circa 1916 as a response to rising prices and labor demand due to the increased number of firms, wage increases failed to keep pace with soaring prices. By the end of the war, the number of labor disputes grew sharply, and the labor movement flourished. Furthermore, the rise of business accelerated urbanization as the rural population moved to cities in search of job opportunities. The share of large cities (with a population of $50,000$ or more) as a percentage of the total population was less than $10$% in 1889 but increased to nearly $16$% in 1920 (Online Appendix Figure A.1a). The growth rate was exceptionally high during WWI, indicating that structural changes through the war accelerated the rise of the urban population. Urbanization has increased public investment in large cities, especially in construction investment in social capital such as roads and water supply systems. [^10] Descriptions of the export-led growths, urbanization, and investments are from Nakamura and Odaka (1989, pp. 21–23) and Nakamura (1971, pp. 142; 147).
Exports in the 1920s were lower because of the appreciation of the yen against the real exchange rate due to the accumulation of domestic and foreign currencies during WWI. As the current account deficit was settled through the disbursement of foreign currencies owned by the government and the Bank of Japan, there was no reduction in the domestic money supply despite the sizeable current account deficit. This resulted in the domestic prices (especially prices of non-tradable goods) remaining high. The first half of the 1920s saw a temporary decline in industrial investment as the capital investment boom of the WWI period came to a halt under declining demand. [^11] Conversely, construction investment by the private sector, including electric power and private railways, increased (Nakamura 1971, p. 145). Descriptions about the post-war economy in this paragraph are from Sawai and Tanimoto (2016, pp. 253–254).
@@
Note: This figure shows the gross national product per capita (solid line) and personal consumption expenditure (dashed line) between 1910 and 1926. The prices are deflated using the CPI in 1934–1935 price.
Source: Created by the author using Ohkawa et. al. (1974, p. 237).
As described, there was a short-term business cycle characterized by the boom caused by WWI and the subsequent recession. However, as a whole, Japan achieved a high economic growth rate relative to other countries in the early 20th century. As shown in Figure 1, the growth rates of personal consumption expenditure were particularly high during and immediately after WWI, with average growth rates reaching 112% in 1915–19 and 118% in 1920–24. This was a high level by international standards. [^12] Total consumption as a percentage of the GNP remained above 80%. Although this level is not apparent among other countries (Kuznets 1962), the growth rate of real personal consumption per capita was relatively high: from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and the rates never became negative except during the Russo-Japanese War period. Descriptions on the personal consumption expenditure are from Nakamura (1971, pp. 2–5; 27; 30–31). This highlights the importance of revealing the consumption behavior of urban households to understand the mechanisms underpinning the rapid Japanese economic growth in this period.
@@ 2.2 Tsukishima: A Representative Manufacturing Area
The population in Tokyo City has been growing since the end of the 19th century, and industrialization due to WWI spurred its growth. Tokyo City’s population share increased from $2.5$% in 1889 to approximately $4$% by 1920.131313In 1920, there were more than $10,000$ municipalities in Japan, suggesting how large Tokyo was. There was a difference of about 1 million between Tokyo and Osaka, the second-largest city by population. Online Appendix Figure A.1a summarizes the trends in the number of people in the top five cities between 1889 and 1920. Throughout the 1910s, the percentage of industrial and commercial workers increased in Tokyo City, building the city as a gigantic commercial and industrial place (Tanimoto 2002, p. 8). Column 1 of Panel A in Table 1 shows that nearly half of the male workers in Tokyo worked in the industrial sector as of 1920.
Figure 2: Administrative Wards in Tokyo City
Notes: The name in each lattice indicates the administrative ward in Tokyo. The gray ward indicates the Kyobashi Ward and an island in the Kyobashi Ward indicates Tsukishima. The border surrounded by Asakusa, Honjyo, Nihonbashi, and Fukagawa wards includes the Sumida River. The borders of each ward are based on the administrative districts in 1920. Geographical coordinate system is based on JGD2000/(B, L).
Source: Created by the author using the official shapefile (Ministry of Land, Infrastructure, Transport and Tourism, database).
The subject of this paper is Tsukishima, an industrial area located southeast of Kyobashi Ward in Tokyo (Figure 2). [^14] The base of the island was reclaimed during the Edo and Meiji periods from a delta at the mouth of the Sumida River. Online Appendix A.3 summarizes the development history of Tsukishima in detail. Column (2) of Panel A in Table 1 shows that Kyobashi had an industrial composition like that of Tokyo City in 1920. Column (3) confirms that Tsukishima is a representative industrial area in the city; approximately $68$% of male workers were employed in the manufacturing sector. In the same year, there were $214$ factories on this island, and approximately $80$% of them belonged to the machinery and related sectors (Department of Health, Ministry of the Interior 1923a, pp 385–389). [^15] According to the census of 1920, the number of workers in the metal and machinery equipment manufacturing industry ($27.1$%) was the highest compared to other industries in the entire city of Tokyo. The same tendency was observed in Kyobashi Ward ($35.5$%). Although similar statistics are not available for Tsukishima, this supports the evidence that Kyobashi was not an area with a unique industrial pattern (Statistics Bureau of the Cabinet 1929c, pp. 85–86; 108–109)). Figure 3 shows the spatial distribution of factories in Tsukishima as of June 30, 1918. Tsukishima had many machine factories and shipyards, and the workers’ residences were clustered there. [^16] For example, a representative shipbuilding plant was located on Ishikawajima, an island in the north of Tsukishima (Figure 3). Ships were strongly influenced by the import substitution that was underway at the beginning of the 20th century, and domestic production of ships progressed (Section 2.1). The production of a hydraulic 80-ton hardened crane by Ishikawajima Shipyard in 1917 is an example of the domestic output of a product that had previously relied on imports (Nakamura and Odaka 1989, pp. 10; 32). The focus of this study is the consumption behavior of these worker households in the industrial sector.
@@
Notes: Blocks in green and light green indicate the shipyards. Blocks in red and pink indicate the factories in the machinery sector. Blocks in blue and light blue indicate the factories in the other manufacturing sectors. Blocks in brown are the warehouses. Small circles colored in black and white show the workplaces in the machinery and other manufacturing sectors, respectively.
Source: Department of Health, Ministry of the Interior 1923c, second map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
@@
Notes:
Panel A: This panel summarizes the industrial structures based on the occupations of male workers measured in the 1920 population census. Each share is calculated as the number of males who worked as regular workers in each sector divided by the total number of male workers (in percentage). All the figures do not include the number of unemployed (mugyō sha) males.
@@
Panel B: The average household size is the number of people divided by the number of households. The share of married males is the number of married males living with/without spouses divided by the total number of males (in percentage). The average age of males is calculated using the tables for population by age group reported in the census. As the tables are open-ended, the class for over 60 years old is rounded as $64.5$ following the range of the former category (i.e., 50–59). All the figures in this panel do not include a small number of quasi-households (jyun setai), which include the person, such as the students in dormitories and the patients in hospitals.
The figures for Tsukishima listed in column 3 of both panels are based on the total sum of statistics of all the blocks (chōme) in Tsukishima (i.e., from Tsukudajima to Tsukishima dōri jyūni chōme).
Sources: Columns 1 and 2 of Panel A use the Statistics Bureau of the Cabinet (1929c, pp. 18–19). Columns 1 and 2 of Panel B use the Statistics Bureau of the Cabinet (1929c, pp. 38–43); Tokyo City Office (1922b, pp. 2–3; 42–46). Column 3 of Panel A uses the Tokyo City Office (1922c, pp. 42–47). Column 3 of Panel B uses the Tokyo City Office (1922a, pp. 26–31); Tokyo City Office (1922b, pp. 2–3; 42–46).
@@
Panel B in Table 1 summarizes the population and occupational statistics from the census to examine the characteristics of working-class households in Tsukishima. First, the average household size in Tsukishima is $4.3$—similar to but slightly smaller than the values in Tokyo City ($4.6$) and Kyobashi Ward ($4.7$). The percentage of the married male population is $40$%, which is similar to the range of the figures for Tokyo ($38$%) and Kyobashi ($36$%). The average age of men in Tsukishima is $25.1$, slightly lower than the figures for Tokyo and Kyobashi. While the sex ratio by age group in Tsukishima shows a similar trend to that of Tokyo and Kyobashi, the ratios for the 1414141419191919 and 2020202029292929 age bins are higher than those. Overall, these statistics reflect the fact that the typical household structure is a couple with two children. The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory worker households. The percentage of married men in Tsukishima suggests that a certain proportion of the households in their early 20s might have been couples before they started raising children.

### v14

# Risk-coping Behaviors in Metropolis: Evidence from Working-class Households in Prewar Tokyo

@@ Abstract
I analyze the risk-coping strategies of factory-worker households in early twentieth-century Tokyo. I digitized a unique daily longitudinal household budget survey conducted in Tsukishima, a representative manufacturing area, to examine how consumption was affected by idiosyncratic shocks. I find that although the households were vulnerable and their consumption levels were impacted by these shocks, the estimated income elasticity of indispensable consumption was relatively low in the short run. The results of the mechanism analysis suggest that credit purchases from local retailers helped smooth short-run consumption, highlighting the role of informal credit institutions in mitigating vulnerability among urban worker households.
@@ 1 Introduction
During the early stages of industrialization, workers who migrated to large cities in search of employment faced various uncertainties. Lacking the communal ties found in rural areas, they were forced to manage on their own. The vulnerability to idiosyncratic income shocks makes it challenging to smooth consumption and increases the risk of falling into poverty. Therefore, understanding how these workers managed their livelihoods helps clarify the mechanisms behind improvements in living standards in industrializing cities.
This study examines risk-coping behaviors among working-class households in Tsukishima, a representative manufacturing area in early twentieth-century Tokyo City. Against the backdrop of industrialization during World War I (WWI), many workers migrated to the city, forming a notable portion of the population. Despite the absence of public insurance systems, evidence shows that these workers were able to form and maintain families during this period (Nakagawa 1985). This historical context offers a rare empirical opportunity to understand the risk-coping strategies of urban workers during industrialization.
To provide micro-level evidence on risk-coping behaviors, I digitize a unique daily longitudinal budget survey of factory-worker households. Using a systematic consumption-smoothing regression, I find that the full insurance hypothesis is strongly rejected for overall consumption expenditure. However, I also find evidence that expenses on indispensable consumption categories are partially smoothed in the short run: the estimated income elasticity for the food category is less than half of that for overall consumption. I then use several temporary income categories within the same regression framework to assess the risk-coping behaviors. Considering both favorable and adverse income shocks, I find that savings, credit purchases, and gifts responded to these shocks. To focus specifically on responses to adverse income shocks, I use the household head’s absence due to the influenza pandemic as an instrument. Results from the instrumental variable approach show that credit purchases from local retailers played a significant role in mitigating adverse idiosyncratic shocks, whereas other risk-mitigation strategies were less practical.
Given these results, I analyze the institutional aspects of credit purchases descriptively. First, I find that credit purchases were common among workers and that volatility in household heads’ earnings did not correlate with average credit purchases. Detailed cases suggest that risk premiums were not charged as explicit interest but were instead embedded in retail prices owing to imperfect customer information. Second, I examine conditions among retailers using several official survey reports. These reports provide suggestive evidence that retailers competed intensely within the same sales areas. Because credit purchases did not require explicit contracts, they often led to losses from unpaid sales. However, to retain regular customers and build loyalty, local retailers had strong incentives to offer credit transactions.
The primary contribution of this study is the revelation of the role of informal credit institutions in a historical metropolis. By considering both formal and informal financial institutions simultaneously, this study documents how informal credit transactions with local retailers helped mitigate idiosyncratic income losses. This finding adds to the literature on consumption-smoothing behavior across various economies, [^1] The early representative literature includes Townsend (1994), Townsend (1995), and Ravallion and Chaudhuri (1997). Morduch (1999), Townsend (1995), and Dercon (2004) review earlier studies. Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions. as well as the studies investigating the mechanisms behind human capital formation through consumption smoothing. [^2] Evidence shows that markets in contemporary developing countries are often imperfect. Consequently, households remain vulnerable to risk, and idiosyncratic shocks affect their consumption levels, influencing human capital formation (Foster 1995; Rose 1999; Gertler and Gruber 2002). Specifically, I present a case study of an urban setting characterized by weak communal ties, which has often been overlooked in the literature. For example, development economics research shows that informal credit institutions based on relatives and neighbors help stabilize consumption in rural economies, especially where formal financial institutions are scarce (Udry 1994, 1990; Fafchamps and Lund 2003). [^3] Related studies examine risk-sharing among households with similar cultural and institutional backgrounds (Grimard 1997; Munshi and Rosenzweig 2016). Recent field experiments also focus on village economies (Comola and Prina 2023; Macours et al. 2022). Townsend (1995), Skoufias and Quisumbing (2005), Aguila et al. (2017), and Ogasawara (2024) provide case studies on middle-sized contemporary and historical economies, but evidence on agglomeration economies such as Tokyo City remains scarce. Another important strand of the literature highlights the role of social insurance provided by religious communities in rural contexts (Ager and Ciccone 2018; Ferrara and Testa 2023; Chen 2010). My findings suggest that, in the case of a metropolis, workers primarily relied on informal credit institutions arising from local commercial practices, as social institutions based on communal ties and religious organizations were largely absent. This result may also offer empirical lessons for rapidly urbanizing cities in developing countries. [^4] The real GDP per capita in 2011 dollars was $3,217$ in 1919 Japan (Bolt and van Zanden 2025). This figure is close to the $3,537$ reported for South and Southeast Asian countries in 2010, where rapid urbanization has been observed (Henderson and Turner 2020).
The second contribution is the incorporation of analytical perspectives on consumption-smoothing behavior into economic history studies. I examine a set of homogeneous households in a representative manufacturing area, which provides a valuable empirical setting for testing consumption smoothing. [^5] Focusing on a sample that shares similar occupational and regional characteristics systematically reduces bias when estimating income elasticity (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). This study advances economic history research that documents compensatory responses to income loss among working-class households but lacks direct evidence on consumption smoothing (Scott and Walker 2012; Horrell and Oxley 2000; Kiesling 1996; Boyer 1997). I implement a systematic panel data analysis to provide comparable estimates of income elasticity. This comparability improves understanding of the availability of risk-coping institutions at different stages of economic development and strengthens the potential for cross-period and cross-economy comparisons. It also adds an analytical perspective to other strands of literature on the supply-side evidence of financial institutions (O’Connell and Reid 2005) and historical patterns of consumption and production (Francks 2009; Gordon 2013).
@@
Lastly, examining how, and to what extent, local society could secure household consumption offers a new perspective within the urban history literature. A representative historical study on urban households shows that workers in Tokyo were able to establish and sustain their families around WWI (Nakagawa 1985). The findings of this study suggest that the expansion of urbanization and credit transactions supports this traditional view. Another implication relates to the findings of Munshi and Rosenzweig (2016), who link spatial wage inequality and men’s mobility in India to the availability of risk-mitigation networks among workers within the same caste. From this perspective, the existence of local credit institutions with lower entry barriers may have encouraged rural workers to migrate to urban areas, helping to explain Tokyo’s rapid growth around WWI.
@@ 2.1 Japanese Economy circa the First World War
Between 1885 and 1915, Japan’s real gross national product (GNP) grew at an annual rate of 2.6%, while the population increased by 1.1%. [^6] Faster economic growth relative to population growth indicates that modern economic growth had begun during this period, as defined by Kuznets (Miyamoto 2008, pp. 60; 64). In 1915, the employed population was distributed across the primary, secondary, and tertiary sectors at 63%, 20%, and 18%, respectively. Based on domestic genuine production, these sectors accounted for just under 30%, over 30%, and approximately 40%, respectively. Descriptions in this subsection are based on Nakamura (1971, pp. 128; 142; 145; 147), Nakamura and Odaka (1989, pp. 10; 21–23), and Sawai and Tanimoto (2016, pp. 250; 253–254). Soon after the outbreak of WWI, European exports declined. Japan increased its exports to Asian markets in place of European suppliers and exported munitions and foodstuffs to European allies, resulting in a current account surplus. Import substitution further expanded the domestic market, driving growth in the machinery, metal, and chemical sectors. A global shortage of ships also spurred rapid development in shipbuilding, which expanded significantly during this period. The growth of the shipbuilding industry, in turn, stimulated demand in the machinery and steel industries. Figure 1 shows that Japan’s nominal GNP grew by 11.511.5% from 1915 to 1919 (4.44.4% in real terms), marking the highest growth rate observed throughout the 1910s and 1920s.
@@
Note: This figure shows Japan’s GNP per capita (solid line) and personal consumption expenditure (dashed line) from 1910 to 1926. Prices are deflated using the 1934–1935 consumer price index (CPI). Source: Created by the author based on Ohkawa et al. (1974, p. 237).
The rise of business accelerated urbanization as the rural population moved to cities in search of employment. The share of large cities (with a population of $50,000$ or more) in the total population was less than 1010% in 1889 but rose to nearly 1616% in 1920 (Online Appendix Figure A.1a). Growth was particularly rapid during WWI, indicating that structural changes during the war accelerated urban population growth. Urbanization also increased public investment in large cities, especially in social infrastructure such as roads and water supply systems.
Exports in the 1920s were lower owing to the appreciation of the yen in real terms, caused by the accumulation of domestic and foreign currencies during WWI. The current account deficit was settled through the disbursement of foreign currencies held by the government and the Bank of Japan, so the domestic money supply remained unchanged despite the sizeable deficit. As a result, domestic prices stayed high. The first half of the 1920s saw a temporary decline in industrial investment as the WWI-era capital investment boom ended under falling demand.
Overall, while a short-term business cycle emerged–marked by the WWI-induced boom and subsequent recession–Japan achieved a high economic growth rate relative to other countries in the early twentieth century. The average growth rates of personal consumption expenditure reached 112% in 1915–1919 and 118% in 1920–1924 (Figure 1), reflecting a high level by international standards. [^7] Total consumption as a percentage of the GNP remained above 80%. Although this level is not internationally high-standard (Kuznets 1962), the growth rate of real personal consumption per capita was relatively high; from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and it was never negative except during the Russo–Japanese War period (Nakamura 1971, pp. 2–5; 27; 30–31).
@@ 2.2 Tsukishima: A Representative Manufacturing Area
Notes: Blocks in green and light green represent shipyards. Blocks in red and pink represent factories in the machinery sector. Blocks in blue and light blue represent factories in other manufacturing sectors. Blocks in brown indicate warehouses. Small circles colored black and white show workplaces in the machinery and other manufacturing sectors, respectively. The upper right corner indicates north. Source: Department of Health, Ministry of the Interior 1923c, second map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
The population of Tokyo City had been growing since the end of the 19th century, and industrialization during WWI further spurred its growth. [^8] Tokyo City’s population share increased from 2.52.5% in 1889 to approximately 44% by 1920. Its population exceeded Osaka’s, the second-largest city, by about 1 million. In 1920, Japan had more than $10,000$ municipalities, underscoring Tokyo’s relative size. Online Appendix Figure A.1a summarizes the population trends of the top five cities between 1889 and 1920. Throughout the 1910s, the share of industrial and commercial workers in Tokyo City increased. According to the 1920 census, nearly half of the male workers in Tokyo were employed in the industrial sector (Column 1 of Panel A in Table 1).
The subject of this study is Tsukishima, an island located southeast of Kyobashi Ward. [^9] Online Appendix Figure A.3 shows the location of Tsukishima. The development history of Tsukishima is summarized in Online Appendix A.3. Kyobashi has an industrial composition similar to that of Tokyo City (Column 2), and Tsukishima represents a typical industrial area within the ward, with approximately 6868% of male workers employed in the manufacturing sector (Column 3). In 1920, Tsukishima had $214$ factories, of which approximately 8080% belonged to the machinery and related sectors. [^10] Department of Health, Ministry of the Interior (1923a, pp. 385–389). According to the 1920 census, the metal and machinery equipment manufacturing industry (27.127.1%) employed the largest share of workers in Tokyo. A similar pattern was observed in Kyobashi Ward (35.535.5%) (Statistics Bureau of the Cabinet 1929c, pp. 85–86; 108–109). Figure 2 shows the spatial distribution of factories in Tsukishima as of June 30, 1918, demonstrating the concentration of machine factories and shipyards. [^11] A representative shipbuilding plant, Ishikawajima Shipyard, was located on Ishikawajima, an island north of Tsukishima. Import substitution during this period strongly accelerated domestic ship production (Section 2.1). The production of a hydraulic 80-ton hardened crane by Ishikawajima Shipyard in 1917 exemplifies domestic output of a product that had previously relied on imports (Nakamura and Odaka 1989, pp. 10; 32).
@@
Notes: Panel A: This panel summarizes the industrial structure based on the occupations of male workers recorded in the 1920 Population Census. Each share is calculated as the number of males employed as regular workers in a given sector divided by the total number of male workers (in percentage). All figures exclude unemployed (mugyō sha) males. Panel B: The average household size is calculated as the total number of people divided by the number of households. The share of married males is the number of married males, living with/without spouses, divided by the total number of males (in percentage). The average age of males is calculated using the population tables by age group reported in the census. For open-ended age categories, the class for those over 60 years old is rounded to $64.5$, following the range of the previous category (50–59). All figures in this panel exclude a small number of quasi-households (jyun setai), which include individuals such as students in dormitories and patients in hospitals. The figures for Tsukishima listed in Column 3 of both panels are based on the total sum of statistics across all blocks (chōme) in Tsukishima (i.e., from Tsukudajima to Tsukishima dōri jyūni chōme). Sources: Columns 1 and 2 of Panel A: Statistics Bureau of the Cabinet (1929c, pp. 18–19). Columns 1 and 2 of Panel B: Statistics Bureau of the Cabinet (1929c, pp. 38–43); Tokyo City Office (1922b, pp. 2–3; 42–46). Column 3 of Panel A: Tokyo City Office (1922c, pp. 42–47). Column 3 of Panel B: Tokyo City Office (1922a, pp. 26–31); Tokyo City Office (1922b, pp. 2–3; 42–46).
@@
Panel B in Table 1 summarizes population and occupational statistics from the census to examine the characteristics of households in Tsukishima. The average household size ($4.3$) is similar to, but slightly smaller than, the values in Tokyo City and Kyobashi Ward. The percentage of married males (4040%) is comparable to the figures in Columns (1) and (2), while the average age of men ($25.1$) is slightly lower. Sex ratios follow a similar trend to those of Tokyo and Kyobashi, although the ratios for the 14141919 and 20202929 age bins are higher. Overall, these statistics suggest that the typical household consists of a couple with two children. The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory-worker households, and the percentage of married men indicates that a portion of households in their early 20s may have been couples before starting a family.